Calelate the pt of buffer thet is moude b mising 65 ml f .20 M NAHCo, and 75 mi of o.16M Nas CO3. mixi
Answer #1

PKa of HCO3^- is 10.33

no of moles of NaHCO3 = molarity * volume in L

                                       = 0.2*0.065   = 0.013 moles

no of moles of Na2Co3   = molarity * volume in L

                                       = 0.16*0.075   = 0.012 moles

PH   = PKa + log[Na2CO3]/[NaHCo3]

        = 10.33 + log0.012/0.013

         = 10.33-0.03476

           = 10.3 >>>>>answer